OpenStudy (anonymous):

Solve the system of equations. 3x - 4y = -10 and 5x + 8y = -2 A. (1, 2) B. (2, 1) C. (1, -2) D. (-2, 1)

OpenStudy (anonymous):

x=-2 y=1

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

You basically plug each of the answer choices in and see if they satisfy each equation: For example: A. (1, 2) that means x=1 and y=2 so you plug in these values for the first equation 3x - 4y = -10 ====> 3(1)-4(2)=-10 ====> 3-8=-10 ===> -5= -10 NOT TRUE! thus (1,2) is not a solution to the first equation, so we eliminate this option. (we don't bother plugging in these values for the second equation because if it doesn't satisfy both it is not a solution) In order for the ordered pair (x,y) to be a solution it must satisfy (must make a true statement) when you plug it into each equation. Another way you can look at it is graphically. The equations both are linear (power of x is 1) so they are lines...basically the question is asking at what point do these equations meet at if they intersect. So that means they share a point in common. A line is made up of an infinite amount of points. For a point (x,y) to be a solution of a linear equation (equation such as y=mx+b where x has power 1) it means the point is on that line.

OpenStudy (zzr0ck3r):

so does maclaurin! but only at the origin

OpenStudy (shane_b):

Or...you could just solve it: \[3x - 4y = -10\]\[5x + 8y = -2\] Take one of the unknown variables and put it in terms of the other unkown: \[x=\frac{4y-10}{3}\] Now plug this in for x in the other equation and solve for y:\[5(\frac{4y-10}{3})+8y=-2\]\[\frac{20y-50}{3}+8y=-2\]\[20y-50+24y=-6\]\[44y-50=-6\]\[44y=44\]\[y=1\] Now that you know y, plug it back into the other equation and solve for x: \[3x-4(1)=-10\]\[3x=-6\]\[x=-2\]

OpenStudy (anonymous):

Yep, there are many ways to solve it...

OpenStudy (zzr0ck3r):

I thought my joke was pretty good...

OpenStudy (anonymous):

oh sorry @zzr0ck3r I was referring to Shane_B...as for your joke HEHEHEHEHE I GET IT :-) AND IT IS FUNNY!!! Derivative is your slope but you find your general derivative equation evaluated at 0...i guess you are a math person too, awesome!!! so Openstudy has a lot I guess. Awesome.

OpenStudy (anonymous):

option 'd' is correct..